# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 2.2. Basic results of large deviations
We recall some basic concepts and results in the theory of large deviations. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $X$ be a topological space with the Borel $\sigma$-field $\mathcal{B}(X)$.
Definition 2.3 (Rate function).
Let $I$ be a function from $X$ to $[0,\infty]$. We say that $I$ is a rate function on $X$ if $I$ is lower semi-continuous on $X$.
Definition 2.4 (Large Derivations).
Let $I$ be a rate function on $X$. Suppose that $\left\{\mu_{\varepsilon}\right\}_{\varepsilon>0}$ is a family of probability measures on $X$. We say that $\left\{\mu_{\varepsilon}\right\}_{\varepsilon>0}$ satisfies the large deviations on $X$ with speed $\varepsilon^{-1}$ and rate function $I$, if
(i) (Upper bound) For any closed subset $F\subset X$,
$\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(F)\leqslant-\inf_{x\in F}I(x).$
(ii) (Lower bound) For any open subset $G\subset X$,
$\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(G)\geqslant-\inf_{x\in G}I(x).$
The following entropy method will be used in our proof of lower bound.
@@
[Mar10, Lemma 7],[GHW24, Lemma 3.2] Let $E$ be a separable and Hausdorff space, $I$ be a function from $E$ to $[0,+\infty],$ and $\{\mu_{\varepsilon}\}_{\varepsilon>0}$ be a family of probability measures on $E$. Then $\{\mu_{\varepsilon}\}$ satisfies the large deviation lower bound with speed $\varepsilon^{-1}$ and rate function $I$ if and only if, for every $x\in E$, there exists a sequence of probability measures $\{\pi_{\varepsilon,x}\}_{\varepsilon>0}$ that satisfies the following two conditions.
@@
2.3. The Regularization of Noise
Recall that $\{e_{k}\}_{k\in\mathbb{Z}^{2}\backslash\{0\}}$ is an orthonormal basis of $H$. A cylindrical Wiener process $W$ on $H$ has the following representation,
where $\{\beta_{k}\}_{k\in\mathbb{Z}^{2}\backslash\{0\}}$ is a family of independent real-valued Brownian motions. For every $\delta>0$ and fixed $s>\alpha+1$, we define the regularizing operator $Q_{\delta}$ as
$Q_{\delta}=(I+\delta(-\Delta)^{s})^{-1},$
then
$\sqrt{Q_{\delta}}e_{k}=(I+\delta|k|^{2s})^{-1/2}e_{k}=:\lambda_{\delta,k}e_{k},\quad k\in\mathbb{Z}^{2}\backslash\{0\}.$
$\|T\|_{HS}^{2}:=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\|Te_{k}\|_{H}^{2}.$ (2.6)
A direct computation shows that
$\left\|\Lambda^{\alpha}\circ\sqrt{Q_{\delta}}\right\|_{HS}^{2}$ $=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}|k|^{2\alpha}\lambda_{\delta,k}^{2}=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\frac{|k|^{2\alpha}}{1+\delta|k|^{2s}}$
This indicates the blow-up speed of solutions of (1.3) and is therefore related to the scaling regimes we will encounter later on.
In the study of large deviations, we restrict that the scaling regime $(\varepsilon,\delta(\varepsilon))$ satisfies
$\lim_{\varepsilon\rightarrow 0}\varepsilon\delta(\varepsilon)^{-\frac{\alpha+1}{s}}=0.$ (2.7)
Similarly, we require that the scaling regime $(\varepsilon,m(\varepsilon))$ for the Galerkin sequence (3.7) satisfies
In this context, we denote $\xi=\mathrm{d}W(t)/\mathrm{d}t$ and $\xi_{\delta}=\sqrt{Q_{\delta}}\mathrm{d}W(t)/\mathrm{d}t.$
@@ 3. Existence and Exponential Tightness
In this section, we will establish the existence of solutions of the regularized stochastic PDE (1.3) for every $\varepsilon,\delta>0$, and the exponential tightness under the scaling regimes (2.7) and (2.8).
First, we provide the definition of solutions. The following definition can be regarded as a generalization of the Leray solution. Since the Sobolev exponent is negative when $\beta=\alpha/2+1/4$, a Littlewood-Paley analysis is needed to make the nonlinear term well-defined.
Definition 3.1.
We say that $(\theta_{\varepsilon,\delta},W)$, $\left(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in[0,T]},\mathbb{P}\right)$ is a stochastic generalized Leray solution of (1.3) with initial data $\theta_{0}$ if, $\theta_{0}$ is an $H^{\alpha-2\beta}(\mathbb{T}^{2})$-valued random element, $W$ is a cylindrical Wiener process on $H^{\alpha-2\beta}(\mathbb{T}^{2})$ independent of $\theta_{0}$, and $\theta_{\varepsilon,\delta}$ is a progressively measurable process satisfying:
(i) $\mathbb{P}$-almost surely, $\theta_{\varepsilon,\delta}\in L^{\infty}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\cap C([0,T];H_{w}^{\alpha-2\beta}(\mathbb{T}^{2}))$.
$\left\langle\theta_{\varepsilon,\delta}(t),\varphi(t)\right\rangle+\int_{0}^{t}\left\langle\theta_{\varepsilon,\delta}(s),\Lambda^{2\alpha}\varphi(s)\right\rangle\,\mathrm{d}s=\left\langle\theta_{0},\varphi(0)\right\rangle+\int_{0}^{t}\langle\theta_{\varepsilon,\delta}(s),\partial_{s}\varphi(s)\rangle\,\mathrm{d}s$
$-\frac{1}{2}\int_{0}^{t}\left\langle R_{2}\theta_{\varepsilon,\delta}(s),\left[\Lambda,\partial_{1}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta}(s)\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\theta_{\varepsilon,\delta}(s),\left[\Lambda,\partial_{2}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta}(s)\right\rangle\,\mathrm{d}s$
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi(s),\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\rangle.$
Here and subsequently, the nonlinear terms $\left\langle R_{j}\theta_{\varepsilon,\delta}(s),\left[\Lambda,\partial_{i}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta}(s)\right\rangle(i,j=1,2)$ are defined by the $H^{-1/2}-H^{1/2}$ dual as in (2.3).
(iii) $\theta_{\varepsilon,\delta}$ satisfies the pathwise $H^{\alpha-2\beta}$-energy inequality: $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
$\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+$ $\int_{0}^{t}\left\|\theta_{\varepsilon,\delta}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{0}\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,\delta}(s),\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\rangle+\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ\sqrt{Q_{\delta}}\right\|_{HS}^{2}t.$ (3.1)
@@
The existence of a martingale solution is established in [ZZ14, Theorem 4.5], following from a similar argument in [FG95]. And the pathwise $H^{\alpha-2\beta}$-energy inequality can be derived in a similar way as in the proof of [GHW24, Proposition 5.1]. We only sketch the proof for the pathwise $H^{\alpha-2\beta}$-energy inequality.
@@
We denote $H_{m}:=\mathrm{span}\{e_{k}\}_{0<|k|\leqslant m}.$ Let $P_{m}$ be the projection operator from $H$ to $H_{m}$. Now we consider the Galerkin approximation equations
$\mathrm{d}\tilde{\theta}_{\varepsilon,\delta,m}(t)=-\Lambda^{2\alpha}\tilde{\theta}_{\varepsilon,\delta,m}(t)\,\mathrm{d}t-P_{m}\left(u_{\tilde{\theta}_{\varepsilon,\delta,m}}(t)\cdot\nabla\tilde{\theta}_{\varepsilon,\delta,m}(t)\right)\,\mathrm{d}t+\varepsilon^{1/2}\Lambda^{2\beta}P_{m}\sqrt{Q_{\delta}}\,\mathrm{d}\tilde{W}(t),$ (3.2)
$\tilde{\theta}_{\varepsilon,\delta,m}(0)=P_{m}\tilde{\theta}_{0}.$
Since the coefficients in (3.2) satisfy the local Lipschitz condition, there exists a local solution on $[0,\tau_{K}]$ for any $K>0$. Here the stopping time $\tau_{K}$ is defined by $\tau_{K}:=\inf\{t\in[0,T]:\|\tilde{\theta}_{\varepsilon,\delta,m}(t)\|_{H}^{2}>K\}\wedge T$ for any $t\in[0,T].$ By the finite dimensional Itô’s formula, $\tilde{\mathbb{P}}$-almost surely, for all $t\in[0,\tau_{K})$,
$\frac{1}{2}\left\|\tilde{\theta}_{\varepsilon,\delta,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $+\int_{0}^{t}\left\|\tilde{\theta}_{\varepsilon,\delta,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\left\|P_{m}\tilde{\theta}_{0}\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ (3.3)
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\tilde{\theta}_{\varepsilon,\delta,m}(s),P_{m}\sqrt{Q_{\delta}}\,\mathrm{d}\tilde{W}(s)\rangle+\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ P_{m}\sqrt{Q_{\delta}}\right\|_{HS}^{2}t.$
Indeed, in the case $\beta=\alpha/2$, since $\nabla\cdot u_{\tilde{\theta}_{\varepsilon,\delta,m}}=0$, it holds that
$\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\cdot\nabla\tilde{\theta}_{\varepsilon,\delta,m},\tilde{\theta}_{\varepsilon,\delta,m}\rangle=-\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\tilde{\theta}_{\varepsilon,\delta,m},\nabla\tilde{\theta}_{\varepsilon,\delta,m}\rangle=0.$
And in the case $\beta=\alpha/2+1/4$, it holds that
$\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\cdot\nabla\tilde{\theta}_{\varepsilon,\delta,m},\Lambda^{-1}\tilde{\theta}_{\varepsilon,\delta,m}\rangle=-\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\tilde{\theta}_{\varepsilon,\delta,m},\nabla\Lambda^{-1}\tilde{\theta}_{\varepsilon,\delta,m}\rangle=-\langle\tilde{\theta}_{\varepsilon,\delta,m}R^{\perp}\tilde{\theta}_{\varepsilon,\delta,m},R\tilde{\theta}_{\varepsilon,\delta,m}\rangle=0.$
Therefore, (3.3) holds. Using Burkholder-Davis-Gundy inequality, we get the estimate
$\mathbb{\tilde{E}}\sup_{t\in[0,\tau_{K})}\|\tilde{\theta}_{\varepsilon,\delta,m}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+2\mathbb{\tilde{E}}\int_{0}^{\tau_{K}}\|\tilde{\theta}_{\varepsilon,\delta,m}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant C(T),$ (3.4)
[…]
$\mathbb{E}$ $\left[\int_{0}^{T}\psi(t)\left(\frac{1}{2}\left\|\theta_{\varepsilon,\delta,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\frac{1}{2}\left\|\theta_{\varepsilon,\delta,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\theta_{\varepsilon,\delta,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right.\right.$ (3.5)
$+\left.\left.\sqrt{\varepsilon}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\,\theta_{\varepsilon,\delta,m}(s),P_{m}\sqrt{Q_{\delta}}\mathrm{d}W(s)\right\rangle-\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ P_{m}\sqrt{Q_{\delta}}\right\|_{HS}^{2}t\right)\,\mathrm{d}t\right]=0.$
Set $M^{m}_{1}(t):=\int_{0}^{t}\sqrt{\varepsilon}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,\delta,m}(s),P_{m}\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\right\rangle$, then $\{M^{m}_{1}\}_{m\geqslant 1}$ is tight in $C([0,T];\mathbb{R}).$ We denote the limit by $M_{1}:$
$\mathbb{E}\left[\int_{0}^{T}\psi(t)M_{1}(t)\,\mathrm{d}t\right]=\lim_{m\rightarrow\infty}\mathbb{E}\left[\int_{0}^{T}\psi(t)M_{1}^{m}(t)\,\mathrm{d}t\right].$
Letting $m\rightarrow\infty$ in (3.5), it follows from the lower semi-continuity of the Sobolev norms and Fatou’s lemma that
$\mathbb{E}\left[\int_{0}^{T}\psi(t)\left(\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right.\right.$ $-\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
$+\int_{0}^{t}\left\|\theta_{\varepsilon,\delta}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ $\left.\left.+M_{1}(t)-\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ P_{m}\sqrt{Q_{\delta}}\right\|_{HS}^{2}t\right)\,\mathrm{d}t\right]\leqslant 0.$
Since $\psi\in L^{\infty}\left(\mathbb{P};L^{\infty}\left([0,T],\mathbb{R}_{+}\right)\right)$ is arbitrary, $\mathbb{P}$-almost surely, for all $t\in[0,T],$
$\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\leqslant\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\int_{0}^{t}\left\|\theta_{\varepsilon,\delta}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s-M_{1}(t)+\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ\sqrt{Q_{\delta}}\right\|_{HS}^{2}t.$
To establish the pathwise $H^{\alpha-2\beta}$-energy inequality (3.1), it is sufficient to verify that
$M_{1}(t)=\sqrt{\varepsilon}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,\delta}(s),\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\right\rangle.$ (3.6)
A detailed proof of (3.6) can be found in [GHW24, Proposition 5.1]. ∎
If we set $\delta=0$ and $\theta_{m,0}=P_{m}\theta_{0}\in H_{m}$, then (3.2) becomes
@@
$\mathrm{d}\theta_{\varepsilon,m}(t)=-\Lambda^{2\alpha}\theta_{\varepsilon,m}\,\mathrm{d}t-P_{m}\left(u_{\theta_{\varepsilon,m}}\cdot\nabla\theta_{\varepsilon,m}\right)\,\mathrm{d}t+\varepsilon^{1/2}\Lambda^{2\beta}P_{m}\,\mathrm{d}W(t),$ (3.7)
$\theta_{\varepsilon,m}(0)=\theta_{m,0}.$
@@
For any $\varepsilon>0$ and $m\in\mathbb{N}_{+}$, (3.7) admits a solution $\theta_{\varepsilon,m}$ as well. For any $\varepsilon,\delta>0$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3). We denote $\mu_{\varepsilon,\delta}$ and $\mu_{\varepsilon,m}$ as the laws of $\theta_{\varepsilon,\delta}$ and $\theta_{\varepsilon,m}$, respectively. In the remainder of this section, we will show that $\{\mu_{\varepsilon,\delta}\}$ and $\{\mu_{\varepsilon,m}\}$ are exponentially tight in $\mathbb{X}_{\alpha,\beta}$ under (2.7) and (2.8) respectively. To prove this, we require some exponential estimate lemmas.
@@
The following lemma is aimed at obtaining an estimate in the space $L^{\infty}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$.
@@
Let $\varepsilon>0$ and $m\in\mathbb{N}$. Suppose that $\theta_{\varepsilon,m}$ is a solution of (3.7) with $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$, then there exists $\eta_{0}>0$ such that for every $\eta\in\left[0,\eta_{0}\right]$, under the scaling regime (2.8), we have
$\varepsilon\log\mathbb{E}\exp\left\{\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\leqslant C\left(\eta,T\right),$
@@
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\left(\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\geqslant R\right)=-\infty.$

### v3

@@ 2.3. Basic results of large deviations
The following entropy method will be used in our proof of the lower bound.
@@
[Mariani10, Lemma 7],[GHW24_LLNS, Lemma 3.2] Let $E$ be a separable and Hausdorff space, $I$ be a function from $E$ to $0,+\infty],$ and $\{\mu_{\varepsilon}\}_{\varepsilon>0}$ be a family of probability measures on $E$. Then $\{\mu_{\varepsilon}\}$ satisfies the large deviation lower bound with speed $\varepsilon^{-1}$ and rate function $I$ if and only if, for every $x\in E$, there exists a sequence of probability measures $\{\pi_{\varepsilon,x}\}_{\varepsilon>0}$ that satisfies the following two conditions.
@@
Let $\mu_{\varepsilon,\delta(\varepsilon)}^{0}:=\mathcal{G}\left(0,\varepsilon Q_{\delta(\varepsilon)}/2\right)$ be the Gaussian measure on $H^{\alpha-2\beta}(\mathbb{T}^{2})$. The following large deviations of initial data satisfying Assumption A is an application of Lemma 2.3, whose conditions are straightforward to verify.
Proposition 2.4.
Assume that the scaling regime (1.7) holds for $\varepsilon,\delta(\varepsilon)>0,$ then $\{\mu_{\varepsilon,\delta(\varepsilon)}^{0}\}_{\varepsilon>0}$ satisfy the large deviations on $H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ with speed $\varepsilon^{-1}$ and rate function $\mathcal{I}_{0}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$.
@@ 3. Existence and Exponential Tightness
In this section, we will establish the existence of solutions of the regularized stochastic PDE (1.3) for every $\varepsilon,\delta(\varepsilon)>0$, and the exponential tightness under the scaling regimes (1.7) and (2.5).
@@
The existence of a martingale solution is established in [ZZ14, Theorem 4.5], following from a standard argument similar to [FG95_Martingale_2dNS]. And the pathwise $H^{\alpha-2\beta}$-energy inequality can be derived in a similar way as in the proof of [GHW24_LLNS, Proposition 5.1]. ∎
@@
For any $\varepsilon>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$, we denote $H_{m(\varepsilon)}:=\mathrm{span}\{e_{k}\}_{0<|k|\leqslant m(\varepsilon)}$ and let $P_{m(\varepsilon)}$ be the projection operator from $H$ to $H_{m(\varepsilon)}$. Setting $\delta=0$ and $\theta_{m(\varepsilon),0}=P_{m(\varepsilon)}\theta_{0}\in H_{m(\varepsilon)}$, the Galerkin approximation equation of (1.3) is given by
@@
$\mathrm{d}\theta_{\varepsilon,m(\varepsilon)}(t)=-\Lambda^{2\alpha}\theta_{\varepsilon,m(\varepsilon)}\,\mathrm{d}t-P_{m(\varepsilon)}\left(u_{\theta_{\varepsilon,m(\varepsilon)}}\cdot\nabla\theta_{\varepsilon,m(\varepsilon)}\right)\,\mathrm{d}t+\varepsilon^{1/2}\Lambda^{2\beta}P_{m(\varepsilon)}\,\mathrm{d}W(t),$ (3.1)
$\theta_{\varepsilon,m(\varepsilon)}(0)=\theta_{m(\varepsilon),0}.$
@@
For any $\varepsilon>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$, (3.1) admits a pathwise unique probabilistically strong solution $\theta_{\varepsilon,m(\varepsilon)}$([RZZ15, Theorem 4.2],[DC06_weak_strong, Theorem 1.1]). For any $\varepsilon,\delta(\varepsilon)>0$, let $\theta_{\varepsilon,\delta(\varepsilon)}$ be a stochastic generalized Leray solution of (1.3). We denote $\mu_{\varepsilon,\delta(\varepsilon)}$ and $\mu_{\varepsilon,m(\varepsilon)}$ as the laws of $\theta_{\varepsilon,\delta(\varepsilon)}$ and $\theta_{\varepsilon,m(\varepsilon)}$, respectively. In the remainder of this section, we will show that $\{\mu_{\varepsilon,\delta(\varepsilon)}\}$ and $\{\mu_{\varepsilon,m(\varepsilon)}\}$ are exponentially tight in $\mathbb{X}_{\alpha,\beta}$ under (1.7) and (2.5) respectively.
@@
The following lemma is aimed at obtaining an estimate in the space $L^{\infty}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$. Based on this estimate, we will derive the estimates in $W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))$ and $C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))$, and then prove the exponential tightness.
@@
Let $\varepsilon>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$. Suppose that $\theta_{\varepsilon,m(\varepsilon)}$ is the solution of (3.1) with $\theta_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)$, then there exists $\eta_{0}>0$ such that for every $\eta\in\left[0,\eta_{0}\right]$, under the scaling regime (2.5), we have
$\varepsilon\log\mathbb{E}\exp\left\{\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\leqslant C\left(\eta,T\right),$
@@
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\big(\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\geqslant R\big)=-\infty.$